BexleyBrooks:

Simplify the expression. Write your answers using integers or improper fractions. 2/3n−2(−2n+ 2/3 )

CherryLover:

uhm idk the answer

jhonyy9:

please more clearly this fraction !!!

buttercat:

. _ .

jhonyy9:

@jhonyy9 wrote:
please more clearly this fraction !!!
understandably

CherryLover:

@jhonyy9 wrote:
please more clearly this fraction !!!
fr fr

jhonyy9:

what is the first term ? (2/3n) ?

buttercat:

whoa you're an honorary professor of mathematics?!

CherryLover:

@jhonyy9 wrote:
what is the first term ? (2/3n) ?
teach me your ways of math!

CherryLover:

@buttercat wrote:
whoa you're an honorary professor of mathematics?!
ikr

jhonyy9:

@cherrylover wrote:
@jhonyy9 wrote:
what is the first term ? (2/3n) ?
teach me your ways of math!
in what problem please ?

jhonyy9:

\[\frac{ 2 }{ 3 }n -2(-2n +\frac{ 2 }{ 3 }) = ?\] this is what you need simplify ?

toga:

@buttercat wrote:
whoa you're an honorary professor of mathematics?!
you need 5,000 medals in math to get that

toga:

i think

AcornTheNut:

this expression does not make sense to anyone,,

toga:

@acornthenut wrote:
this expression does not make sense to anyone,,
nope I tried to figure it out and couldn't.

adrianluvvsyouu2:

Can you please send an image of the problem or write it using LaTeX (if you know how)?

marienne:

the answer is \(\frac{2}{3}(7n-2)\)

BexleyBrooks:

@jhonyy9 wrote:
what is the first term ? (2/3n) ?
2 over 3n

BexleyBrooks:

@jhonyy9 wrote:
\[\frac{ 2 }{ 3 }n -2(-2n +\frac{ 2 }{ 3 }) = ?\] this is what you need simplify ?
yes

BexleyBrooks:

@adrianluvvsyouu2 wrote:
Can you please send an image of the problem or write it using LaTeX (if you know how)?
not sure how to do that

DemonKinge:

W pfp

BexleyBrooks:

@buttercat wrote:
whoa you're an honorary professor of mathematics?!
what?

BexleyBrooks:

Ya'll I figured out the answer X'D